# Adapting Multi-modal Large Language Model to Concept Drift in the Long-tailed Open World

### v1

@@ 3.1 Taming the Tailed Drift and OOD Drift for Robust Fine-tuning
We compare our proposed vision-language model with other specialized models to explicitly demonstrate its superior performance in long-tailed open-world scenarios. As shown in Table 1, our model demonstrates exceptional overall performance in long-tailed classification across all three large-scale datasets. The results validate the robustness of our vision language model against biases arising from tail drift, particularly when leveraging large-scale data for both pre-training and fine-tuning.
Table 1: Evaluation results of long-tailed classification on ImageNet-LT, iNatualist2018 and Places-LT datasets with Vit-Large. The best-performing models are highlighted in red. Many, Medi. and Few denote the evaluated splits of many-shot (>100 training samples), medium-shot (20-100 samples) and few-shot (<20 samples). Top-1 accuracy is applied to evaluate the performance of different methods.
Methods ImageNet-LT iNaturalist 2018 Places-LT
Many Medi. Few All Many Medi. Few All Many Medi. Few All
MiSLAS [26] 62.0 49.1 32.8 51.4 - - - 70.7 - - - 38.3
Bl-Softmax [27] 64.1 48.2 33.4 52.3 - - - 70.6 - - - 39.4
LADE [28] 64.4 47.7 34.3 52.3 - - - 69.3 - - - 39.2
ACE [29] - - - 56.6 - - - 72.9 - - - -
RIDE [30] 68.2 53.8 36.0 56.9 70.9 72.4 73.1 72.6 - - - -
PaCo [31] 68.2 58.7 41.0 60.0 70.3 73.2 73.6 73.2 36.1 47.9 35.3 41.2
NCL [32] - - - 60.5 - - - 74.9 - - - 41.8
MEM [33] 76.0 62.0 43.7 64.9 83.6 83.0 81.3 82.4 49.1 47.4 37.4 46.0
Ours 80.3 71.1 37.2 65.6 84.3 83.1 80.7 82.6 61.6 53.3 32.6 46.9
Furthermore, our model exhibits superior performance on the Many and Medium splits, while it slightly lags behind other models in the Few split. On one hand, our findings align with the scaling law observed in large models, where the performance exhibits a power law relationship with the size of the data. As the amount of data increases, the model’s feature extraction capability becomes more robust leading to improved performance. On the other hand, our goal is to alleviate the bias induced by tail drift, instead of averaging the feature space of each category and sacrificing head categories to enhance the performance of tail categories.
@@
Methods SVHN LSUN iSUN Texture Average
FPR AUROC FPR AUROC FPR AUROC FPR AUROC FPR AUROC
ProxyAnchor [34] 87.2 82.4 37.2 91.7 70.0 85.0 65.6 85.0 65.0 86.0
CE + SimCLR [35] 24.8 94.5 56.4 89.0 66.5 83.8 63.7 82.0 52.9 87.3
CSI [36] 44.5 92.7 75.6 83.8 76.6 85.0 61.6 86.5 64.6 87.0
SSD+ [37] 31.2 94.2 79.4 85.2 80.9 84.1 66.6 86.2 64.5 87.4
KNN+ [11] 39.2 92.8 49.0 89.3 75.0 82.7 57.2 88.4 55.1 88.3
CIDER [19] 12.6 97.8 30.2 92.8 46.0 88.9 35.6 92.3 31.1 93.0
Ours 8.3 98.7 20.3 97.5 32.5 95.2 45.1 96.3 26.6 96.9
@@
Table 3: Evaluation results of image-text alignment of different contrastive learning strategies in the stage of pre-training, from three perspectives: ID intra-class compactness, ID inter-class separability and the separability between ID and OOD categories. The cosine metric is utilized to measure these distances, which is expressed as average degrees with standard deviation in brackets. We compare our proposed tSP with classical cosine loss, under balanced scenario (BL, ImageNet) and imbalanced scenarios (LT, ImageNet-LT), respectively.
@@
LT / tSP 36.2 ($\pm$3.53) 85.6 ($\pm$1.88) 101.3 ($\pm$1.25)
Moreover, we verified at the feature-level that the proposed T-distributed adapter significantly alleviates the bias from tailed drift and OOD drift in the pre-training. As exhibited in Table 3, the degree of ID intra-class compactness reduces from 49.2 in LT/cosine to 36.2 in LT/tSP. It thereby validates the effectiveness of the proposed T-distributed adapter in enhancing feature extraction in long-tailed scenarios. Notably, the decrease in standard deviation demonstrates that the model considerably mitigates the bias induced by tail drift. In addition, our method achieves remarkable inter-class separability within in-distribution categories under long-tailed scenarios, even surpassing the performance of cosine achieved under the balanced dataset. It confirms the effectiveness of the proposed high kurtosis method in enhancing the alignment between images and text in the pre-training stage. Concerning the separability between ID and OOD, we achieve superior results even than the balanced condition, attributed to the inherent light-tailed property of the T-distributed adapter. It ensures our approach performs robust for OOD drift detection based on the spherical distribution robustly.
@@ 3.3.2 Various Concentration Parameter $\kappa$
Acc 61.3 63.8 59.9 57.2 62.3
In Table 5, setting the concentration parameter to $\kappa=16$ yields superior results. We argue that, a smaller concentration makes it challenging to effectively mitigate the biases introduced by tail drift and OOD drift in the vision language model. In terms of the bigger concentration, the model is hard to train due to the high kurtosis of the tSP metric. In the context of concentration as a trainable parameter with the initialization of 16, there is a slight reduction in model performance, accompanied by a marginal increase of the concentration parameter to 16.37. We assert that, the introduction of the new parameter increases the model’s complexity, thereby making the training process more challenging. Additionally, the low learning rate hinders the effective training of the concentration parameters, leading its less change compared with the initialization.

### v2

@@ 3.1 Taming the Tailed Drift and OOD Drift for Robust Fine-tuning
We compare our proposed vision-language model with other models to explicitly demonstrate its superior performance in long-tailed open-world scenarios. As shown in Table 1, our model demonstrates exceptional overall performance in long-tailed classification across two large-scale datasets, namely the ImageNet-LT and iNaturalist 2018. Especially, training from scratch means that our method uses the imbalanced dataset for pre-training instead of utilizing the pre-trained model, such as clip [23] pre-trained by the large WIT dataset. The results validate the robustness of our vision language model against biases arising from tailed drift, particularly when leveraging large-scale data for both pre-training and fine-tuning.
Table 1: Evaluation results of long-tailed classification on ImageNet-LT and iNatualist2018. The best-performing models are highlighted in red. Many, Medi. and Few denote the evaluated splits of many-shot ($>$100 training samples), medium-shot (20-100 samples) and few-shot ($<$20 samples). Top-1 accuracy is applied to evaluate the performance of different methods.
ImageNet-LT iNaturalist 2018
Methods Backbones Many Medium Few All Many Medium Few All
Training from scratch
cRT [24] 61.8 46.2 27.3 49.6 69.0 66.0 63.2 65.2
LWS [24] 60.2 47.2 30.3 49.9 65.0 66.3 65.5 65.9
MiSLAS [25] 62.9 50.7 34.3 52.7 73.2 72.4 70.4 71.6
BALMS [26] 64.1 48.2 33.4 52.3 - - - 70.6
LADE [27] 64.4 47.7 34.3 52.3 - - - 69.3
ACE [28] - - - 56.6 - - - 72.9
RIDE [29] 68.2 53.8 36.0 56.9 70.9 72.4 73.1 72.6
PaCo [30] 68.2 58.7 41.0 60.0 70.3 73.2 73.6 73.2
NCL [31] ResNet-50 - - - 57.4 72.0 74.9 73.8 74.2
ViT [9] 73.7 46.5 15.6 52.4 65.4 55.3 50.9 54.6
MAE [32] 74.7 48.2 19.4 54.5 79.6 70.8 65.0 69.4
DeiT [33] 70.4 40.9 12.8 48.4 72.9 62.8 55.8 61.0
LiVT [34] ViT-B/16 73.6 56.4 41.0 60.9 78.9 76.5 74.8 76.1
Ours 77.2 68.6 51.3 69.6 83.3 82.1 80.5 81.5
Only Fine-tuning
LPT [35] - - - - 62.1 76.2 79.3 76.1
BALLAD [36] 79.1 74.5 69.8 75.7 - - - -
Decoder [37] - - - 73.2 - - - 59.2
VL-LTR [38] 84.5 74.6 59.3 77.2 - - - 81.0
LIFT [39] ViT-B/16 (w/ Pretrained Clip) 80.2 76.1 71.5 77.0 72.4 79.0 81.1 79.1
Ours 79.9 76.8 74.5 77.6 84.1 82.7 81.0 82.1
Compared to other methods trained from scratch, especially the ViT model, our model demonstrates a notable lead across all metrics, indicating our effective mitigation of concept drift during the pre-training, and providing robust pre-trained models for downstream tasks. Furthermore, to compare the current long-tailed methods, we the same setup as LIFT[39], i.e., using the pre-trained model of clip and only fine-tuning. The superior results on medium and few splits of ImageNet-LT demonstrate the adaptability and robustness of our model in dealing with the gradual drift caused by tail data. Besides, although we are slightly behind LIFT[39] in the few split of iNatualist2018, we still surpass it overall, exhibiting that our method does not compromise the accuracy of the head category to improve the tailed.
@@
Methods SVHN LSUN iSUN Texture
FPR AUROC FPR AUROC FPR AUROC FPR AUROC
ProxyAnchor [40] 87.2 82.4 37.2 91.7 70.0 85.0 65.6 85.0
CE + SimCLR [41] 24.8 94.5 56.4 89.0 66.5 83.8 63.7 82.0
CSI [42] 44.5 92.7 75.6 83.8 76.6 85.0 61.6 86.5
SSD+ [43] 31.2 94.2 79.4 85.2 80.9 84.1 66.6 86.2
KNN+ [8] 39.2 92.8 49.0 89.3 75.0 82.7 57.2 88.4
CIDER [16] 12.6 97.8 30.2 92.8 46.0 88.9 35.6 92.3
Ours 8.3 98.7 20.3 97.5 32.5 95.2 45.1 96.3
@@ 3.2 Concept Drift-Aware Image-Text Alignment for Effective Pre-training
Table 3: Evaluation results of image-text alignment of different contrastive learning strategies in the stage of pre-training, from three perspectives: ID intra-class compactness, ID inter-class separability and the separability between ID and OOD categories. The cosine metric is utilized to measure these distances, which is expressed as average degrees with standard deviation in brackets. We compare our proposed Thp with classical cosine loss, under balanced scenario (BL, ImageNet) and imbalanced scenarios (LT, ImageNet-LT), respectively.
@@
LT / Thp 36.2 ($\pm$3.53) 85.6 ($\pm$1.88) 101.3 ($\pm$1.25)
Moreover, we verified at the feature-level that the proposed T-distributed adapter significantly alleviates the bias from tailed drift and OOD drift in the pre-training. As exhibited in Table 3, the degree of ID intra-class compactness reduces from 49.2 in LT/cosine to 36.2 in LT/Thp. It thereby validates the effectiveness of the proposed T-distributed adapter in enhancing feature extraction in long-tailed scenarios. Notably, the decrease in standard deviation demonstrates that the model considerably mitigates the bias induced by tail drift. In addition, our method achieves remarkable inter-class separability within in-distribution categories under long-tailed scenarios, even surpassing the performance of cosine achieved under the balanced dataset. It confirms the effectiveness of the proposed high kurtosis method in enhancing the alignment between images and text in the pre-training stage. Concerning the separability between ID and OOD, we achieve superior results even than the balanced condition, attributed to the inherent light-tailed property of the T-distributed adapter. It ensures our approach performs robust for OOD drift detection based on the spherical distribution robustly.
@@ 3.3.2 Various Concentration Parameter $\kappa$
Acc 67.2 69.6 64.9 61.1 68.3
In Table 5, setting the concentration parameter to $\kappa=16$ yields superior results. We argue that, a smaller concentration makes it challenging to effectively mitigate the biases introduced by tail drift and OOD drift in the vision language model. In terms of the bigger concentration, the model is hard to train due to the high kurtosis of the Thp metric. In the context of concentration as a trainable parameter with the initialization of 16, there is a slight reduction in model performance, accompanied by a marginal increase of the concentration parameter to 16.37. We assert that, the introduction of the new parameter increases the model’s complexity, thereby making the training process more challenging. Additionally, the low learning rate hinders the effective training of the concentration parameters, leading its less change compared with the initialization.